Find the equation of a line that passes through the point (2,1) and has a gradient of 2.

Leave your answer in the form y

=

m

x

+

c



1
Expert's answer
2021-12-06T17:10:11-0500

Let us find the equation of a line that passes through the point (2,1) and has a gradient of 2.

Leave our answer in the form "y\n\n=\n\nm\n\nx\n\n+\n\nc".

Let us use the formula "y=y_0+m(x-x_0)." In our case, "x_0=2,y_0=1,m=2."

It follows that "y=1+2(x-2)=1+2x-4=2x-3."

therefore, the equation of the line is "y=2x-3."